find inverse laplace transorm : se^-2s/(s^2 + pi^2)





1
Expert's answer
2022-01-24T18:51:09-0500

"\qquad\text{By property of Laplace Transform,}\\
if\ \mathcal{L}^{-1}(G(s))=g(t), then\\
\mathcal{L}^{-1}(e^{-as}G(s))=u(t-a)\times g(t-a)\\
\quad\\
Now,\mathcal{L}(f(t))=\frac{se^{-2s}}{s^2+\pi^2}=F(s)\\
\Rightarrow f(t)=\mathcal{L}^{-1}(F(s))=\mathcal{L}^{-1}(\frac{se^{-2s}}{s^2+\pi^2})=\mathcal{L}^{-1}(e^{-2s}\times \frac{s}{s^2+\pi^2})\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad=\mathcal{L}^{-1}(e^{-2s}\times G(s)), where\ G(s)=\frac{s}{s^2+\pi^2}\\
But, \mathcal{L}^{-1}(G(s))=\mathcal{L}^{-1}(\frac{s}{s^2+\pi^2})=\cos(\pi t)=g(t), thus\\
\Rightarrow f(t)=\mathcal{L}^{-1}(e^{-2s}\times G(s))\\
\qquad\quad =u(t-2)\times g(t-2), \text{from the above stated property where a=2}\\
\qquad\quad =u(t-2)\times \cos(\pi(t-2)), since\ g(t)=\cos(\pi t)\\
Thus, \mathcal{L}^{-1}(\frac{se^{-2s}}{s^2+\pi^2})=u(t-2)\times \cos(\pi(t-2))."